A wooden cykindrical pole is 7 m high and its base radius is 10 cm. Find its weight if the wood weighs 225 kg per cubic metre.
Answer by Cromlix(4381)

Volume = Pi * r^2 * h
Volume = Pi * 0.1^2 * 7
Volume = 0.22 m^3 (2 decimal places)
Weight = 0.22 * 225
Weight = 49.5 kg
Hope this helps.
:-)